OpenStudy (anonymous):

how about if it is fraction? wat will I do 1st? 3x/4 ≥-6 Thank you!

OpenStudy (anonymous):

1. Multiply both sides by 4 2. Divide both sides by 3

OpenStudy (anonymous):

Or 1. Multiply both sides by 4/3

OpenStudy (anonymous):

ok, thank you

OpenStudy (anonymous):

i can't remember if it is negative when you subtract a lower positive number from a higher negative number? Help pls. Thank you!

OpenStudy (anonymous):

2-3 = -1

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

yes. a big heLp.... now I don't have hesitation in my answer! Thanks

OpenStudy (anonymous):

wat if negative divide by negative? is it equal to positive?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

+ve times +ve = +ve +ve times -ve = -ve -ve times -ve = +ve Same holds for division.

OpenStudy (anonymous):

ahm...okay, I thought It was different rule!

OpenStudy (anonymous):

I got 2 unfinished answers w/ my questions, bcoz now I'm confused which answer will i proceed, which is correct 3x/4 times 4/3 is equal to 12x/7 or 9x*16?

OpenStudy (anonymous):

\[\frac{3x}{4} ≥-6\]\[\frac{4}{3}\times\frac{3x}{4} ≥-6\times \frac{4}{3}\]Simplify the left and right side.

OpenStudy (anonymous):

I have 2 different answers bcoz I can't remember if it's correct to cross multiply it!

OpenStudy (anonymous):

\[\frac{4}{3}\times\frac{3x}{4} ≥-6\times \frac{4}{3}\]\[\frac{4\times3x}{3\times 4} ≥\frac{-6\times 4}{3}\]

OpenStudy (anonymous):

If you don't understand / can't handle this, just use the first method I told you: 1. Multiply both sides by 4 2. Divide both sides by 3

OpenStudy (anonymous):

I've got x≥-8 is my answer correct?

OpenStudy (help!!!!):

-1

OpenStudy (anonymous):

-8 is right...

OpenStudy (anonymous):

I mean x≥-8

OpenStudy (anonymous):